Try this» Object thrown vertically upwards with ini. Vel. U .covers equal distance in 7th &8th sec. What is displacemet in 8th sec?? (g=10

Try this»
Object thrown vertically upwards with ini. Vel. U .covers equal distance in 7th &8th sec. What is displacemet in 8th sec?? (g=10

Grade:12

3 Answers

naresh masaboina
10 years ago

320m

Simran Bhatia
10 years ago

Displacement in the n-th second = u + (a/2) * (2n -1)

It is increasing monotonically with n.

So the only way distance travelled in 7th second = distance traveled in 8th second is when at t=7sec the object is at its maximum height.

In which case, displacement in 7th second = (-1) * displacement in 8th second

Or, u - 5 * (2*7 -1) = - [ u - 5 * (2*8 -1)]

Which gives u = 70 m/s

So displacement in 8th second = | 70 - 5 * (2*8 -1) | =  | 70 - 75| = 5 meters

suru siva
10 years ago

the time of ascent must be 7 seconds. 

u=10*7 = 70

h max = 245

distance travelled in 8th second=5m

so displacement after 8 seconds is 240 m